OpenStudy (blondey_315):

A 68-mile trip uses two gallons of gas. How many gallons would you use on a 110-mile trip? Round to the nearest tenth of a gallon. 1.2 3.2 3.8 4.1

OpenStudy (mathstudent55):

Use a proportion.

OpenStudy (blondey_315):

how would I set it up?

OpenStudy (anonymous):

\[\frac{68}{2}=\frac{110}{x}\] or \[\frac{2}{68}=\frac{x}{110}\] either way

OpenStudy (mathstudent55):

First, write two fraction lines: 

OpenStudy (mathstudent55):

Choose the numerators to be miles, and the denominators to be gallons. In each fraction you write a corresponding set of miles and gallons.